OpenStudy (carolina4567):

A segment with endpoints C (3, 4) and D (11, 3) is divided by a point E such that CE and DE form a 3:5 ratio. Find the x value for E. 6 5.5 7.5 7

OpenStudy (carolina4567):

@John_ES

OpenStudy (carolina4567):

HELP Just 4 More Questions Left Please Help

OpenStudy (carolina4567):

Im thinking C

OpenStudy (carolina4567):

HELP

OpenStudy (john_es):

I think is also C. But I found 7.8, instead 7.5.

OpenStudy (carolina4567):

ok how about this one Find the perimeter of a quadrilateral with vertices at C (−1, 1), D (3, 4), E (6, 0), F (2, −3). Round your answer to the nearest hundredth when necessary. 12 units 16 units 20 units 24 units

OpenStudy (carolina4567):

Im thinking B

OpenStudy (john_es):

The right is C.

OpenStudy (carolina4567):

Use the following four coordinates to determine which one has a distance of 6 units from point E (3, 1). (5, 1) (2, 4) (3, 7) (7, 2)

OpenStudy (carolina4567):

I would say D

OpenStudy (john_es):

It is C, again.

OpenStudy (carolina4567):

Last One Point X is located at (2, 4), and Y is located at (−6, 10). Find the y value of the point that lies halfway between X and Y. 2 7 -3 -8

OpenStudy (carolina4567):

@John_ES

OpenStudy (carolina4567):

Im thinking C

OpenStudy (john_es):

It is B. the mid point can be found with, \[\frac{A+B}{2}=\frac{(2,4)+(-6,10)}{2}=(-2,7)\]

OpenStudy (carolina4567):

55/60 THANKS

OpenStudy (john_es):

You're welcome.